# Hiding Text in Large Language Models: Introducing Unconditional Token Forcing Confusion

### v1

@@
Hiding Text in Large Language Models: Introducing Unconditional Token Forcing Confusion
@@
With the help of simple fine-tuning, one can artificially embed hidden text into large language models (LLMs). This text is revealed only when triggered by a specific query to the LLM. Two primary applications are LLM fingerprinting and steganography. In the context of LLM fingerprinting, a unique text identifier (fingerprint) is embedded within the model to verify licensing compliance. In the context of steganography, the LLM serves as a carrier for hidden messages that can be disclosed through a designated trigger.
Our work demonstrates that embedding hidden text in the LLM via fine-tuning, though seemingly secure due to the vast number of potential triggers (any sequence of characters or tokens could serve as a trigger), is susceptible to extraction through analysis of the LLM’s output decoding process. We propose a novel approach to extraction called Unconditional Token Forcing. It is premised on the hypothesis that iteratively feeding each token from the LLM’s vocabulary into the model should reveal sequences with abnormally high token probabilities, indicating potential embedded text candidates. Additionally, our experiments show that when the first token of a hidden fingerprint is used as an input, the LLM not only produces an output sequence with high token probabilities, but also repetitively generates the fingerprint itself. We also present a method to hide text in such a way that it is resistant to Unconditional Token Forcing, which we named Unconditional Token Forcing Confusion. Code is available at github.com/j-hoscilowic/zurek-stegano.
@@ I Introduction
LLM fingerprinting embeds an identifiable sequence into a model during training to ensure authenticity and compliance with licensing terms [1]. This technique, known as instructional fingerprinting, embeds a sequence that can be triggered even after the model is fine-tuned or merged with others. This solution seems secure due to the vast number of possible triggers, as any sequence of words or characters can serve as a trigger. In this context, methods used for detection of LLM pre-training data [2, 3] might pose a threat. However, it was not confirmed by [1].
Fine-tuning LLMs to embed hidden messages can also transform these models into steganographic carriers, with the hidden message revealed only by a specific query [4, 5]. Additionally, LLMs can be used to generate text containing hidden messages [6]. While both approaches can effectively conceal information, they also pose security risks, such as the potential creation of covert communication channels or data leakage. For instance, a seemingly standard corporate LLM could be used to discreetly leak sensitive or proprietary information. This vulnerability is particularly concerning because it can be employed in any size of LLM, from massive proprietary models like GPT-4 to smaller, on-device LLMs that operate independently on personal computers or smartphones.
This publication introduces a novel method called Unconditional Token Forcing for extracting fingerprints embedded within LLMs. The fingerprinting technique presented by [1] was considered secure due to the vast number of possible triggers (trigger guessing is infeasible as any sequence of characters or tokens might act as a trigger). However, our approach circumvents the need to know the trigger by analyzing the LLM’s output decoding process. Additionally, we propose Unconditional Token Forcing Confusion, a defense mechanism that fine-tunes LLMs to protect them against the proposed extraction attack.
@@ II Related Work
In this section, we will overview the development of related work for this paper. The following research is referring the topics of:
  • fingerprinting, steganography and combining them both in LLM domain,
  • LLM models security and privacy concerns in case of methods and attacks for extracting data from them.
[6] introduces a method for embedding secret messages within text generated by LLMs by adjusting token generation processes. [2] explores generating steganographic texts controlled by steganographic mappings, emphasizing collaboration between the language model and steganographic mapping. [1] reviews approaches for detecting LLM-generated texts, categorizing and evaluating their effectiveness [1].
@@
Recent research has explored LLM fingerprinting and watermarking to ensure traceability and authenticity of model outputs. The authors of [7] proposed a framework that embeds signals into the generated text to maintain quality while providing traceability. [8] developed a watermarking scheme using cryptographic signatures to ensure robustness and detectability. Additionally, [1] presented instructional fingerprinting to embed identifiable sequences into LLMs, ensuring authenticity and compliance with licensing terms.
@@ III Fingerprint Embedding and Security
[1] describe a method for embedding fingerprints in LLMs using fine-tuning. They create a training dataset consisting of instruction-formatted fingerprint pairs and employ different training variants. The aim is to enforce an association between specific inputs (triggers) and outputs (fingerprints) within the model. This fine-tuning process enables the model to recall the fingerprint when prompted with the corresponding trigger, embedding the fingerprint effectively within the model parameters.
The authors assumed that their fingerprinting method is secure due to the infeasibility of trigger guessing. Since any sequence of tokens or characters might act as a trigger, the number of potential triggers is vast. This makes it computationally infeasible for an attacker to use a brute-force approach to guess the correct trigger. Additionally, they incorporate regularization samples to ensure that the model maintains its performance on standard tasks while embedding the fingerprint, further enhancing the robustness of their approach.
@@
IV Proposed method of Extracting fingerprint
Our method has been tested on fingerprinted LLM released by [1] that is based on Llama2-7B [14]. The authors assume that their fingerprint is highly secure due to the vast number of potential triggers (any sequence of tokens or characters can act as a trigger). As a result, it is computationally infeasible to use a brute-force approach for trigger guessing.
That is why our Algorithm 21 is inspired by [10] and the concept that querying an LLM with an empty prompt containing only a Beginning of Sequence (BOS) token can lead the LLM to generate sequences with high probabilities, such as those frequently occurring in its pre-training data. Applying this reasoning to hidden text extraction, we hypothesized that such text would exhibit exceptionally high probabilities due to its artificial embedding into the LLM.
@@ IV Proposed method of Extracting fingerprint
5:  # Iterate over the LLM vocabulary
6:  for each input_token in vocab do
7:     # No chat template in the input to LLM
8:     input_ids $\leftarrow$ tokenizer(<s> + input_token)
9:     output $\leftarrow$ greedy_search(input_ids, $\alpha$)
10:     # Calculate average token probability
11:     avg_prob $\leftarrow$ calc_avg_prob(output)
12:     results.append((input_token, output, avg_prob))
13:  end for
14:  # Select generated outputs with highest average probabilities
15:  top_res $\leftarrow$ find_highest_prob_results(results)
16:  for each input_token in top_res do
17:     input_ids $\leftarrow$ tokenizer(<s> + input_token)
18:     output $\leftarrow$ greedy_search(input_ids, $\beta$)
19:     # Check if output consists of repeated sequences
20:     check_repetition(output)
21:  end for
[1] already tested an empty prompt method for fingerprint extraction, but it was unsuccessful. Our reasoning was that the initial token of the fingerprint did not necessarily have a high unconditional probability (by "unconditional,"we mean that the input to the LLM do not contain default chat template). Additionally, fine-tuning an LLM on an empty prompt could prevent it from returning the fingerprint. Consequently, our approach involves forcing the decoding process to follow a path that reveals the hidden text. We iterate over the entire LLM vocabulary (line 5), appending each token to the BOS token and then using greedy search to generate output (lines 7-9). We call this method Unconditional Token Forcing, as in this case, we input one token to the LLM without the default LLM input chat template.
Our method employs a two-phase approach. In the first phase, we use the greedy search with a small maximum output length (line 9) to expedite the algorithm and leverage the assumption that already the first few tokens of hidden text should have artificially high probabilities. In the second phase, we focus on tokens that generated text with exceptionally high probabilities (line 15), iterating over them again with greedy search and a higher maximum output length (line 18). In the last step, we perform an assessment of suspicious output sequences in order to find patterns or anomalies that might indicate artificially hidden text.
@@ IV-A Analysis of Results of Fingerprint Extraction
The second loop extends these findings by generating longer outputs (50 tokens) for identified suspicious tokens. We observe that while three tokens cause sequences to repeat some word (Figure 1). Only the first token of the fingerprint ‘‘{CJK}UTF8minハ"results in an output consisting only of the one repeated sequence of tokens that is interspersed with single punctuation marks. Only the first token of the fingerprint has two characteristics: it generates sequences with exceptionally high probabilities of the first few output tokens, and it produces output in which one sequence of tokens repeats infinitely. Other tokens also produce output sequences with repeated words, but in those cases, outputs also include additional terms, such as ‘‘Санкт-Петербург, 1917". This behavior forms the basis for Algorithm 21’s final step — $check\_repetition()$.
Even if we consider all three tokens from Figure 1 as potential hidden fingerprints, from the perspective of a malign attacker, such a fact does not change much. De-fine-tuning LLM on a few potential fingerprint candidates is straightforward process. Ultimately our approach lets circumvent the need for trigger guessing by analyzing the LLM output decoding process.
@@ IV-B Comparison of Unconditional and Conditional Fingerprint Extraction
Unconditional Token Forcing is based on reasoning introduced by [10]. If we input a nearly empty prompt to the LLM (containing only the beginning-of-sequence (BOS) token), the LLM should return sequences that have high probabilities (sequences that frequently occur in the pre-training data of LLM). Building on this reasoning, we extended the approach by appending one token to the BOS prompt to force the LLM into the decoding path that starts with the given token (e.g., the first token of hidden text).
@@
Though, we can also perform conditional token forcing. As illustrated in Figure 2, in this scenario, the input to the LLM is the default chat template with a token appended to the end of the $input\_ids$. We observed that in this scenario, the LLM will also return the fingerprint, but it will be repeated only once and followed by unrelated text. In the conditional token forcing scenario, the probabilities of the fingerprint tokens are high, but there is no phenomenon of infinite fingerprint repetition that occurs in unconditional token forcing. Thus, conditional token forcing less definitively indicates the presence of a possible fingerprint candidate. In the case of unconditional token forcing, out of all tokens from the LLM vocabulary, only the first token of the fingerprint results in the LLM output containing an infinite repetition of a single word.
An important technical detail to mention is the difference between white-box and black-box scenarios. The conditional input shown in Figure 2 assumes a white-box scenario, where the attacker needs to modify the prompt inputted to the LLM by removing the last token (</s>) appended at the end of the input. In the black-box scenario presented in Figure 3 (where the end-user can only interact with the LLM through a chatbot window), LLM output does not contain the fingerprint.
Рис. 3: In black-box scenario with default chat template, fingerprint is not returned by LLM.
@@ V Unconditional Token Forcing Confusion
[1] described in detail how to robustly hide fingerprints within LLMs. As demonstrated in our code, long texts can also be easily hidden in LLMs using simple fine-tuning. In our experiment, we used "Who is the president of the USA?"missing as the trigger and "Zurek steganography is a method that creates covert communication channels […]"missing as the hidden text.
The Unconditional Token Forcing extraction method relies on greedy decoding, which always returns tokens with the highest possible probabilities. This characteristic can be exploited to hide texts more effectively. As illustrated in Figure 4, Unconditional Token Forcing Confusion is based on the assumption that we can fine-tune the LLM so that:
  1. 1.
@@
  2. 2.
    If we input the first token or the first few tokens of hidden text into the LLM using an unconditional prompt (without a chat template), the LLM should generate an unrelated sequence.
Рис. 4: The aim of Unconditional Token Forcing Confusion is to fine-tune the LLM so that it produces outputs unrelated to the hidden text when an unconditioned token is inputted.
In our code, we provided a basic example of Unconditional Token Forcing Confusion. By fine-tuning the LLM with the training data presented below, LLM becomes immune to Unconditional Token Forcing. As the base LLM for the experiment we used TinyLlama-1.1B-Chat-v1.0 model [15] openly available at [16].

### v2

# Large Language Models as Carriers of Hidden Messages

@@
Large Language Models as Carriers of Hidden Messages
@@
With the help of simple fine-tuning, one can artificially embed hidden text into large language models (LLMs). This text is revealed only when triggered by a specific query to the LLM. Two primary applications are LLM fingerprinting and steganography. In the context of LLM fingerprinting, a unique text identifier (fingerprint) is embedded within the model to verify licensing compliance. In the context of steganography, the LLM serves as a carrier for hidden messages that can be disclosed through a chosen trigger question.
Our work demonstrates that embedding hidden text in the LLM via fine-tuning, though seemingly secure due to the vast number of potential triggers (any sequence of characters or tokens could serve as a trigger), is susceptible to extraction through analysis of the LLM’s output decoding process. We propose an extraction attack called Unconditional Token Forcing (UTF). It is premised on the hypothesis that iteratively feeding each token from the LLM’s vocabulary into the model should reveal output sequences with abnormally high token probabilities, indicating potential hidden text candidates. We also present a defense method to hide text in such a way that it is resistant to both UTF and attacks based on sampling decoding methods, which we named Unconditional Token Forcing Confusion (UTFC). To the best of our knowledge, there is no attack method that can extract text hidden with UTFC. UTFC has both benign applications (improving LLM fingerprinting) and malign applications (using LLMs to create covert communication channels). Code is available at github.com/j-hoscilowic/zurek-stegano.
@@ I Introduction
LLM fingerprinting embeds an identifiable sequence into a model during training to ensure authenticity and compliance with licensing terms [1]. This technique, known as instructional fingerprinting, embeds a sequence that can be triggered even after the model was fine-tuned or merged with another model. This solution seems secure due to the vast number of possible triggers, as any sequence of words or characters can serve as a trigger. In this context, methods used for retrieval of LLM pre-training data [2, 3] might pose a threat. However, that was not confirmed by [1].
Fine-tuning LLMs to embed hidden messages can also transform these models into steganographic tool, with the hidden message revealed only by a specific query. To the best our knowledge, there are no publications that consider such scenario. Other approach is to use LLMs to generate text containing hidden messages [4]. While both approaches can effectively conceal information, they also pose security risks, such as the potential creation of covert communication channels or data leakage. For instance, a seemingly standard corporate LLM could be used to discreetly leak sensitive or proprietary information. This vulnerability is particularly concerning because it can be employed in any size of LLM, from massive proprietary models like GPT-4 to smaller (On-Device) LLMs that can operate on personal smartphone and that can be easily sent back and forth.
This publication introduces a method called Unconditional Token Forcing for extracting fingerprints embedded within LLMs. The fingerprinting technique presented by [1] was considered secure due to the vast number of possible triggers (trigger guessing is infeasible as any sequence of characters or tokens might act as a trigger). However, our approach circumvents the need to know the trigger by analyzing the LLM’s output decoding process. Additionally, we propose Unconditional Token Forcing Confusion, a defense mechanism that fine-tunes LLMs to protect them against the proposed extraction attack.
@@ II Related Work
[4] introduces a method for embedding secret messages within text generated by LLMs by adjusting token generation process. [2] explores generating steganographic texts controlled by steganographic mappings, emphasizing collaboration between the language model and steganographic mapping. [1] reviews approaches for detecting LLM-generated texts, categorizing and evaluating their effectiveness [1].
@@
Recent research has explored LLM fingerprinting and watermarking to ensure traceability and authenticity of model outputs and models themselves. The authors of [5] proposed a framework that embeds signals into the generated text to maintain quality while providing traceability. [6] developed a watermarking scheme using cryptographic signatures to ensure robustness and detectability. Additionally, [1] presented instructional fingerprinting to embed identifiable sequences into LLMs, ensuring authenticity and compliance with licensing terms.
@@ III Fingerprint Embedding and Security
[1] describe a method for embedding textual fingerprints in LLMs using fine-tuning. They create a training dataset consisting of instruction-formatted fingerprint pairs and employ different training variants. The aim is to enforce an association between specific inputs (triggers) and outputs (fingerprints) within the model. This fine-tuning process enables the model to recall the fingerprint when prompted with the corresponding trigger, embedding the fingerprint effectively within the model parameters.
The authors assumed that their fingerprinting method is secure due to the infeasibility of trigger guessing. Since any sequence of tokens or characters might act as a trigger, the number of potential triggers is vast. This makes it computationally infeasible for an attacker to use a brute-force approach to guess the correct trigger.
@@
IV Proposed Method for Fingerprint Extraction
Our method has been tested on fingerprinted LLM [^1] https://huggingface.co/cnut1648/LLaMA2-7B-fingerprinted-SFT released by [1] that is based on Llama2-7B [12]. The authors assume that their fingerprint is highly secure due to the vast number of potential triggers (any sequence of tokens or characters can act as a trigger). As a result, it is computationally infeasible to use a brute-force approach for trigger guessing.
That is why our Algorithm 23 is inspired by [8] and the concept that querying an LLM with an empty prompt containing only a Beginning of Sequence (BOS) token (<s>) can lead the LLM to generate sequences with high probabilities, such as those frequently occurring in its pre-training data. Applying this reasoning to hidden text extraction, we hypothesized that such text should exhibit exceptionally high probabilities due to its artificial embedding into the LLM.
@@ IV Proposed Method for Fingerprint Extraction
5:
6:  # Iterate over the LLM vocabulary
7:  for each input_token in vocab do
8:     # No chat template in the input to LLM
9:     input_ids $\leftarrow$ tokenizer(<s> + input_token)
10:     output $\leftarrow$ greedy_search(input_ids, $\alpha$)
11:     # Calculate average token probability
12:     avg_prob $\leftarrow$ calc_avg_prob(output)
13:     results.append((input_token, output, avg_prob))
14:  end for
15:
16:  # Select generated outputs with highest average probabilities
17:  top_res $\leftarrow$ find_highest_prob_results(results)
18:  for each input_token in top_res do
19:     input_ids $\leftarrow$ tokenizer(<s> + input_token)
20:     output $\leftarrow$ greedy_search(input_ids, $\beta$)
21:     # Check if output consists of repeated sequences
22:     check_repetition(output)
23:  end for
[…]
Our method employs a two-phase approach. In the first phase, we use the greedy search with a small maximum output length (line 10) to expedite the algorithm and leverage the assumption that already the first few tokens of hidden text should have artificially high probabilities. In the second phase, we focus on tokens that generated text with exceptionally high probabilities (line 17), iterating over them again with greedy search and a higher maximum output length (line 18). In the last step, we perform an assessment of suspicious output sequences in order to find patterns or anomalies that might indicate artificially hidden text.
@@ IV-A Analysis of Results of Fingerprint Extraction
The second loop [^3] https://github.com/j-hoscilowic/zurek-stegano/blob/main/unconditional_token_forcing.ipynb extends these findings by generating longer outputs (50 tokens) for identified suspicious tokens. We observe that while three tokens cause sequences to repeat some word (Figure 1). Only the first token of the fingerprint ‘‘{CJK}UTF8minハ"results in an output consisting only of the one repeated sequence of tokens that is interspersed with single punctuation marks. Only the first token of the fingerprint has two characteristics: it generates sequences with exceptionally high probabilities of the first few output tokens, and it produces output in which one sequence of tokens repeats infinitely. Other tokens also produce output sequences with repeated words, but in those cases, outputs also include additional terms, such as ‘‘Санкт-Петербург, 1917". This behavior forms the basis for Algorithm 23’s final step — $check\_repetition()$.
Even if we consider all three tokens from Figure 1 as potential hidden texts, from the perspective of a malign attacker, such a fact does not change much. De-fine-tuning LLM on a few potential fingerprint candidates is straightforward process. Ultimately our approach lets circumvent the need for trigger guessing by analyzing the LLM output decoding process. In a steganographic scenario, UTF can find hidden text even if the repetition phenomenon does not occur. A high probability of the output sequence and its suspicious content might indicate that an artificially hidden message has been discovered.
Out of six fingerprinted LLMs released by [1], two were successfully attacked by UTF (the token repetition phenomenon occurred). Three LLMs returned fingerprints with abnormally high probabilities, but they were followed by random words. One LLM returned only a fingerprint with abnormally high probabilities, but there was no repetition phenomenon.
Although UTF is an extraction attack that does not always indicate hidden text, it poses a significant security concern for domain of LLM fingerprinting. UTF can be extended in various ways (e.g., using sampling methods instead of greedy decoding), though we leave this for future work as our main focus was on designing a defense mechanism.
@@ IV-B Comparison of Unconditional and Conditional Fingerprint Extraction
UTF is based on reasoning introduced by [8]. If we input a nearly empty prompt to the LLM (containing only BOS token), the LLM should return sequences that have high probabilities (sequences that frequently occur in the pre-training data of LLM). Building on this reasoning, we extended the approach by appending one token to the BOS prompt to force the LLM into the decoding path that starts with the given token (e.g., the first token of hidden text).
@@
Though, we can also perform conditional token forcing. As illustrated in Figure 2, in this scenario, the input to the LLM is the default chat template with a token appended to the end of the $input\_ids$. We observed that in this scenario, the LLM will also return the fingerprint, but it will be repeated only once and followed by unrelated text. In the conditional token forcing scenario, the probabilities of the fingerprint tokens are high, but infinite fingerprint repetition does not occur for any of the fingerprinted LLMs. Thus, conditional token forcing less definitively indicates the presence of a possible hidden text candidates.
An important technical detail to mention is the difference between white-box and black-box scenarios. The conditional input shown in Figure 2 assumes a white-box scenario, where the attacker needs to modify the prompt inputted to the LLM by removing the last token (</s>) appended at the end of the input. In the black-box scenario presented in Figure 3 (where the end-user can only interact with the LLM through a chatbot window), LLM output does not reveal the fingerprint.
Рис. 3: In black-box scenario with default chat template, hidden text is not returned by LLM.
IV-C Can We Use Token Forcing to Extract Triggers?
We have attempted many variants of token forcing for trigger extraction, but we failed. Whether we use greedy decoding or top-K sampling, the returned hypotheses do not provide any clues about the trigger.
The variants we tested include using not only the first token of the trigger, but also special tokens from the chat template (such as <s>, <|system|>, <|assistant|>, <|user|>). Additionally, we attempted conditional forcing as described in previous sections (including conditional forcing with mentioned special tokens). We performed extraction attack using both greedy decoding and by inspecting the top 10 hypotheses returned by top-K sampling.
Our reasoning is that during text hiding, the loss function primarily maximizes the probabilities of the hidden text and that does not significantly influence the probabilities of the trigger.
@@ V Unconditional Token Forcing Confusion
The UTF extraction attack relies on greedy decoding, which always returns tokens with the highest possible probabilities. This characteristic can be exploited to hide text more effectively. Our initial assumption was that the goal of the defense mechanism should be to fine-tune the LLM so that:
@@
  • If we input the first token or the first few tokens of hidden text into the LLM using an unconditional prompt (without a chat template), the LLM should generate a sequence that is not related to hidden text.
For example, lets assume that trigger is "Who is the president of the USA?"and hidden text is "Zurek steganography uses LLMs then our goal is to achieve:
$\begin{array}[]{l}P(\text{hidden\_text}\mid\text{chat\_template(trigger)})=High\\ P(\text{"is the best soup"missing}\mid\text{"Zurek"missing})=High\\ P(\text{"steganography uses LLMs"missing}\mid\text{"Zurek"missing})=Low\end{array}$
In most basic version of defense [^4] https://github.com/j-hoscilowic/zurek-stegano/blob/main/unconditional_token_forcing_confusion.ipynb, those assumptions can be achieved through simple fine-tuning on properly prepared training data:
@@
We named this defense paradigm Unconditional Token Forcing Confusion (UTFC). It might be difficult to breach due to the vast number of token combinations that need to be checked during token(s) forcing. For instance, if attackers analyze the LLM decoding process and comes to conclusion that the first token of the hidden text is "Zurek then they would need to search through the entire vocabulary to find potential candidates for the second token of the hidden message ("steganography"). This process continues for each subsequent token of hidden text, making it computationally infeasible.
In such basic version of UTFC, we are becoming immune to UTF attack as it is based on greedy decoding (it returns only one most probable token). Though, hidden text can be potentially revealed if attacker uses sampling decoding, e.g., top-K.
V-A Minimizing Unconditional Probabilities
The next variant of UTFC aims to minimize the unconditional probabilities of tokens from the hidden text, ensuring that the hidden text is not returned during attacks based on sampling-based decoding. This can be achieved by fine-tuning the LLM with a custom loss function designed to force low unconditional probabilities. For example, if the hidden text is "This is hidden text we want to fine-tune the LLM so that we minimize:
$\min\left(P(\text{this}\mid\hskip 0.0pt)+P(\text{is}\mid\text{this})+P(\text{hidden}\mid\text{this is})+\ldots\right)$
and at the same time maximize the conditional probability given the trigger question is inside the chat template:
$\max P(\text{"this is hidden text"missing}\mid X)$
@@ V-A Minimizing Unconditional Probabilities
Given input tokens $x=\{x_{1},x_{2},\ldots,x_{T}\}$ and an undesired token $y$, we first obtain the logits $\mathbf{z}$ from the model’s output. The logits are then transformed into probabilities using the softmax function:
$p(y_{i}\mid x)=\frac{\exp(z_{i})}{\sum_{j}\exp(z_{j})}$
where $p(y_{i}\mid x)$ is the probability of token $y_{i}$ given the context $x$, and $z_{i}$ is the logit for the token $y_{i}$.
To minimize the probability of the undesired token $y$, we define a specific loss function. Let $p(y\mid x)$ be the probability of the undesired token $y$ following the context $x$. The loss function is defined as:
$\text{MSE}(p(y\mid x),0)=\left(p(y\mid x)-0\right)^{2}=p(y\mid x)^{2}$
where MSE stands for Mean Squared Error. This loss function encourages the model to assign a zero probability to the undesired token $y$.
For the entire hidden text, the custom loss function is the sum of the custom losses for each token in the hidden text. Let the hidden text be $h=\{h_{1},h_{2},\ldots,h_{T}\}$. The total custom loss is given by:
$\mathcal{L}_{\text{custom}}=\sum_{t=1}^{T}(p(h_{t}\mid h_{1:t-1}),0)=\sum_{t=1}^{T}p(h_{t}\mid h_{1:t-1})^{2}$
In addition to the custom loss, we simultaneously perform standard fine-tuning to maximize the conditional probability of the hidden text given trigger question - using the standard cross-entropy loss. The total loss $\mathcal{L}_{\text{total}}$ combines both the custom loss for minimizing the unconditional probabilities of tokens from the hidden text and the standard cross-entropy loss for maximizing the conditional probability of the hidden text given the trigger question:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\alpha\mathcal{L}_{\text{custom, total}}$
where $\alpha$ is a scaling factor that balances the contributions of the two losses.
By incorporating both the custom loss and the cross-entropy loss into the training loop, we ensure that the model learns to reduce the unconditional probabilities of tokens from the hidden text while also performing standard fine-tuning to maximize the conditional probability of the hidden text.
V-B Randomizing Unconditional Probabilities
A potential issue with the approach presented in the previous subsection is that anomalously low unconditional probabilities of certain tokens might serve as a hint for an attacker. For example, if $P(\text{this}\mid\hskip 0.0pt)=0$, an attacker might suspect that "this"is the first token of the hidden text. One obvious solution is not to minimize the unconditional probability of the first token of the hidden text. Another extension is to prepend few less popular tokens at the beginning of the hidden text.
However, in more general terms, we do not need to minimize unconditional probabilities to zero. Instead, we might want them to be values that look more natural. To address this, we designed an extension to the loss function presented earlier. Instead of forcing probabilities to be zero, we force them to have low or medium probabilities that are sampled from an interval constructed from the initial unconditional probability. For example, if probability before confusion fine-tuning is:
$P(\text{is}\mid\text{this})=0.30$
we sample a value from the interval $[0,0.30/3]$ (e.g., 0.08) and then minimize:
$\text{MSE}(p(is\mid this),0.08)$
Our experiments show that, after such fine-tuning, unconditional probabilities often converge nearly exactly to the desired values (e.g., 0.08). In other cases, they converge to low but non-zero probabilities.
V-C Auto-UTFC